OpenStudy (anonymous):

A fruit dealer bought 36 pineapples. He sold some at $1.60 each and the rest at $1.80 each, thereby receiving $61.60. How many did he sell at each price?

OpenStudy (anonymous):

will u help me on my prob.???????

OpenStudy (anonymous):

16 at 1.6 and 20 at 1.8

OpenStudy (anonymous):

Working required??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

let th no of pines sold at 1.60 = x then th no of pines sold at 1.60 = (36-x)

OpenStudy (anonymous):

Money earned from x pines sold at 1.60 = 1.60x Money earned from (36-x) pines sold at 1.80 = 1.80 (36-x) = 64.8 - 1.80x

OpenStudy (anonymous):

ATQ, 1.60x + (64.8 - 1.80x) = 61.60 or 1.60x -1.80x = 61.60 - 64.8 - 0.2x = - 3.2 x = -3.2/ -0.2 x = 16

OpenStudy (anonymous):

so no of pines sold at 1.6 = x =16 and no of pines sold at 1.8 = 36 -16 = 20

OpenStudy (anonymous):

Thks Har

OpenStudy (anonymous):

most welcome